# : Generalizing Large Language Models for E-commerce

### v1

@@ B.2 Instructions on Product Relation Prediction ()
B.3 Instructions on Product Matching ()
@@
B.4 Instructions on Sentiment Analysis ()
Seed Instruction Given the user's review, identify the user's sentiment from the listed options. Answer using one of the options.
Generated Instruction 1 Assess the user's sentiment in the provided review and select the appropriate sentiment option from the list as the answer.
@@ B.4 Instructions on Sentiment Analysis ()
Generated Instruction 3 Carefully assess the user's review for any strong expressions of sentiment, either positive or negative. Based on your analysis, select the most fitting sentiment option from the provided list as output.
Generated Instruction 4 Analyze the user's review text and determine the overall sentiment expressed, then choose the corresponding sentiment option from the provided list (A: very positive, B: positive, C: neutral, D: negative, E: very negative) based on the identified sentiment.
Unseen Instruction Analyze the user's review and determine the sentiment based on the listed options.
B.5 Instructions on Sequential Recommendation ()
@@
Generated Instruction 2 Rank the items in options and predict the user's next purchase from the listed options by analyzing her historical purchases.
Generated Instruction 3 The user's purchase history implies her preferences. Rank the items in the options based on the userβs preferences. Output the item that the user is most likely to purchase next from the options.
@@
Unseen Instruction Estimate the user's intent based on the user's purchase history, and predict the next product that the user is most likely to purchase from the given options.
B.6 Instructions on Multi-class Product Classification ()
@@
B.7 Instructions on Product Substitute Identification ()
@@
Generated Instruction 3 Check if a product can function as a substitute for a given query, even if it doesn't fully meet all requirements. Output yes if it can or no otherwise.
@@
B.8 Instructions on Query Product Ranking ()
@@
Unseen Instruction Evaluate the query against each product's title, determine the relevance between the query and the product, and organize the products in descending order of relevance, ensuring that the product with the highest relevance is positioned at the top of the list.
B.9 Instructions on Answerability Prediction ()
@@
B.10 Instructions on Answer Generation ()
@@
Appendix C Training and Evaluation Details
The training details of general-purpose and e-commerce baseline LLMs are shown in TableΒ A3. The general-purpose LLMs undergo the 1-shot evaluation, which measures the sample with one in-context example and one test case. For each task, the 1-shot evaluation dataset is composed of all 1K test samples and the same amount of training samples (except for , which uses all 253 test samples and 253 training samples) from its own test and training sets. The e-commerce LLM undergoes both 1-shot and 0-shot evaluation.
β
General-purpose LLMs |
URL |
Accessibility |
β
GPT-4 Turbo |
API |
β
Gemini Pro |
API |
β
Claude 2.1 |
https://docs.anthropic.com/claude/reference/getting-started-with-the-api |
API |
β
Llama-2 13B-chat |
Checkpoint |
β
Mistral-7B-Instruct-v0.2 |
Checkpoint |
β
EcomGPT |
Checkpoint |
Tasks |
Baseline Models |
Model Source |
Mode |
Parameters |
SUOpenTagΒ (Xu etΒ al., 2019) |
training |
epoch: 100 |
AVEQAΒ (Wang etΒ al., 2020) |
training |
epoch: 89; batch size: 16 |
RGCNΒ (Schlichtkrull etΒ al., 2018) |
training |
epochs: 500K |
DeBERTaV3Β (He etΒ al., 2022) |
fine-tuning |
epoch: 3; Β Β Β Β batch size: 8 |
| BERTweetΒ (Nguyen etΒ al., 2020) |
https://huggingface.co/finiteautomata/bertweet-base-sentiment-analysis |
training |
epoch: 3; Β Β Β Β batch size: 8 |
DeBERTaV3Β (He etΒ al., 2022) |
fine-tuning |
epoch: 3; Β Β Β Β batch size: 8 |
P5Β (Geng etΒ al., 2022) |
evaluate on checkpoint |
checkpoint: toy base |
gSASRecΒ (Petrov & Macdonald, 2023) |
training |
sequence length: 50 |
RecformerΒ (Li etΒ al., 2023) |
fine-tuning |
epoch: 16; Β Β Β Β batch size 8 |
GPT-4 TurboΒ (OpenAI, 2023) |
1-shot Β Β Β learning |
prompt template inΒ D.2 |
BERTΒ (Devlin etΒ al., 2019) |
fine-tuning |
epoch: 3; Β Β Β Β batch size: 8 |
DeBERTaV3Β (He etΒ al., 2022) |
fine-tuning |
epoch: 3; Β Β Β Β batch size: 8 |
@@ Appendix C Training and Evaluation Details
In this table, ``training" means training from scratch, ``evaluate on checkpoint" means that we evaluate using the checkpoint provided by the link, ``fine-tuning" means that we fine-tune the checkpoint provided by the link on specific tasks, and ``1-shot learning" indicates that we directly use the model checkpoint from the link, and prompt the model with one in-context example.
The training details of SoTA task-specific models are presented in Table A4. For each task, we train and test on the SoTA task-specific models of each task using its own training, validation, and IND test sets (i.e., task-specific). For the tasks with the OOD test set, we save the trained model and test them on the OOD test set. For SUOpenTagΒ (Xu etΒ al., 2019) and AVEQAΒ (Wang etΒ al., 2020), we evaluate the task according to the equationΒ 1. Note that we do not consider MAVEQAΒ (Yang etΒ al., 2022) as the baseline of , as it has demonstrated similar performance to AVEQA. For RGCNΒ (Schlichtkrull etΒ al., 2018), we construct a product-product graph, in which the nodes are products, and the edges are specified by the relations between products (e.g., similar products). We initialize the node embeddings in the RGCN using the embeddings of product titles generated from BERTΒ (Devlin etΒ al., 2019). The BERT, DeBERTa, and BERTweet baselines are implemented through sentence-transformerΒ (Reimers & Gurevych, 2019). For P5Β (Geng etΒ al., 2022), we conduct zero-shot evaluations on beauty, sports, and toys base checkpoints, and report the best result (toy base). Similar to the RGCN, we use the BERT-generated embeddings of product titles to initialize item embeddings in the gSASRec. For both gSASRec and RecformerΒ (Li etΒ al., 2023), we evaluate the results of next product-of-interest within the candidate list as detailed in AppendixΒ A.2.2
Appendix D Prompt Templates
The following prompts are used for evaluating general-purpose and e-commerce LLMs. For Claude 2.1, the prompt is begin with {HUMAN PROMPT} and end with {AI PROMPT}. For Mistral-7B Instruct-v0.2, the prompt is wrapped with ``[INST]".
D.1 Prompt Templates for Zero-shot Evaluation
@@
D.2 Prompt Templates for One-shot Evaluation
@@
Appendix E Analysis on Base Models
@@
In this table, ``Base Model" presents the base models used in models. On each task, the best performance of , , and when using different base models is in bold.
@@
We compare the 6 base models considered for the series of models (i.e., , , and ) and show the performance comparison in TableΒ A5. From TableΒ A5, we observe that Llama-2 13B-chat is the best-performing base model for . The instruction-tuned Llama-2 13B-chat model demonstrates considerable improvement compared to instruction-tuned Flan-T5 XXL on 7 out of the 10 tasks. We also observe that Mistral-7B Instruct-v0.2 and Phi-2 are the best-performing base model for and , respectively. Particularly, with instruction tuning, Mistral-7B Instruct-v0.2 achieves a notable average improvement of 2.2% over Llama-2 7B-chat across the 10 tasks. Similarly, Instruction-tuned Phi-2 also outperforms Flan-T5 XL on 6 out of the 10 tasks, and achieves a similar performance with Flan-T5 XL on the rest 4 tasks. The variations of trainable size and focused aspect contribute to the distinct inherent capabilities of the base models, which play a crucial role in adapting LLMs to e-commerce scenarios.
Appendix F Complete Experimental Results
@@
The complete results for both IND and OOD evaluations for tasks (1) attribute value extraction (), (2) product relation prediction (), (3) product matching (), (4) sentiment analysis (), (5) sequential recommendation (), (6) multi-class product classification (), (7) product substitute identification (), (8) query product ranking (), (9) answerability prediction () and (10) answer generation () are presented in TableΒ A6, Β A7, Β A8, Β A9, Β A14, Β A10, Β A11, Β A14, Β A12, andΒ A13,respectively. Overall, these tables show that models fine-tuned on outperform the general-purpose LLMs and SoTA task-specific models in the IND test. Meanwhile, the exhibits good generalizability to the OOD data. Because of the high-quality , achieves remarkable performance with different base models. Note that #failed in tables represents the number of failure cases that we cannot extract meaningful results from the model output. These failure cases are counted as wrong predictions when calculating the metrics. NPR refers to the negative prediction rate.
| Model | IND | OOD | ||||||||
| Recall* | Precision* | F1* | #failed | Recall* | Precision* | F1* | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.422 | 0.598 | 0.495 | 6 | 0.317 | 0.529 | 0.397 | 1 | |
| Gemini Pro | 0.318 | 0.523 | 0.396 | 4 | 0.203 | 0.426 | 0.275 | 6 | ||
| Claude 2.1 | 0.310 | 0.494 | 0.381 | 59 | 0.312 | 0.600 | 0.410 | 66 | ||
| Llama-2 13B-chat | 0.002 | 0.002 | 0.002 | 0 | 0.000 | 0.000 | 0.000 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.321 | 0.435 | 0.369 | 69 | 0.217 | 0.337 | 0.264 | 52 | ||
| E-commerce LLM | EcomGPT | 0.000 | 0.000 | 0.000 | 905 | 0.001 | 0.042 | 0.001 | 869 | |
| SoTA task-specific model | SUOpenTag | 0.603 | 0.500 | 0.546 | 0 | 0.124 | 0.173 | 0.144 | 0 | |
| AVEQA | 0.425 | 0.491 | 0.456 | 0 | 0.283 | 0.257 | 0.269 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.298 | 0.519 | 0.378 | 7 | 0.362 | 0.701 | 0.477 | 0 | |
| Llama-2 13B-chat | 0.544 | 0.666 | 0.599 | 3 | 0.448 | 0.613 | 0.518 | 2 | ||
| Llama-2 7B-chat | 0.531 | 0.660 | 0.588 | 1 | 0.323 | 0.499 | 0.392 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.720 | 0.799 | 0.757 | 5 | 0.374 | 0.544 | 0.443 | 0 | ||
| Flan-T5 XL | 0.258 | 0.449 | 0.328 | 7 | 0.276 | 0.538 | 0.365 | 3 | ||
| Phi-2 | 0.304 | 0.570 | 0.397 | 265 | 0.288 | 0.488 | 0.362 | 0 | ||
| Generalist | Flan-T5 XXL | 0.353 | 0.611 | 0.447 | 2 | 0.360 | 0.699 | 0.476 | 0 | |
| Llama-2 13B-chat | 0.530 | 0.646 | 0.582 | 1 | 0.276 | 0.425 | 0.335 | 0 | ||
| Llama-2 7B-chat | 0.514 | 0.641 | 0.571 | 4 | 0.236 | 0.392 | 0.294 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.612 | 0.722 | 0.662 | 0 | 0.304 | 0.463 | 0.367 | 0 | ||
| Flan-T5 XL | 0.297 | 0.514 | 0.376 | 1 | 0.267 | 0.518 | 0.352 | 1 | ||
| Phi-2 | 0.455 | 0.578 | 0.509 | 0 | 0.237 | 0.417 | 0.302 | 0 | ||
@@
In this table, ``IND" and ``OOD" indicates in-domain evaluation and out-of-domain evaluation, respectively; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together. Recall*, Precision* and F1* are defined as equationΒ 1 in AppendixΒ A.1.1, and #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. On each task, the best performance is in bold.
| Model | IND | OOD | ||||||||||
| Accuracy | M-Rec | M-Pre | M-F1 | #failed | Accuracy | M-Rec | M-Pre | M-F1 | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.384 | 0.487 | 0.381 | 0.326 | 0 | 0.488 | 0.496 | 0.392 | 0.392 | 0 | |
| Gemini Pro | 0.128 | 0.385 | 0.352 | 0.136 | 1 | 0.147 | 0.359 | 0.390 | 0.123 | 0 | ||
| Claude 2.1 | 0.508 | 0.347 | 0.344 | 0.275 | 10 | 0.362 | 0.394 | 0.400 | 0.277 | 4 | ||
| Llama-2 13B-chat | 0.473 | 0.333 | 0.333 | 0.333 | 0 | 0.419 | 0.338 | 0.339 | 0.324 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.442 | 0.323 | 0.325 | 0.324 | 0 | 0.422 | 0.338 | 0.351 | 0.327 | 0 | ||
| E-commerce LLM | EcomGPT | 0.147 | 0.101 | 0.101 | 0.091 | 444 | 0.125 | 0.125 | 0.092 | 0.096 | 455 | |
| SoTA task-specific model | DeBERTaV3 | 0.762 | 0.575 | 0.620 | 0.588 | 0 | 0.658 | 0.514 | 0.570 | 0.507 | 0 | |
| RGCN | 0.615 | 0.665 | 0.637 | 0.506 | 0 | 0.576 | 0.373 | 0.372 | 0.356 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.754 | 0.516 | 0.511 | 0.508 | 0 | 0.663 | 0.506 | 0.468 | 0.466 | 0 | |
| Llama-2 13B-chat | 0.769 | 0.530 | 0.517 | 0.521 | 0 | 0.690 | 0.520 | 0.472 | 0.483 | 0 | ||
| Llama-2 7B-chat | 0.774 | 0.541 | 0.628 | 0.537 | 0 | 0.695 | 0.526 | 0.803 | 0.498 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.782 | 0.547 | 0.689 | 0.543 | 0 | 0.711 | 0.532 | 0.808 | 0.502 | 0 | ||
| Flan-T5 XL | 0.704 | 0.467 | 0.496 | 0.460 | 0 | 0.592 | 0.471 | 0.625 | 0.427 | 0 | ||
| Phi-2 | 0.584 | 0.372 | 0.379 | 0.348 | 0 | 0.406 | 0.349 | 0.334 | 0.251 | 0 | ||
| Generalist | Flan-T5 XXL | 0.769 | 0.531 | 0.517 | 0.522 | 0 | 0.703 | 0.533 | 0.648 | 0.499 | 0 | |
| Llama-2 13B-chat | 0.775 | 0.599 | 0.635 | 0.611 | 0 | 0.726 | 0.564 | 0.611 | 0.558 | 0 | ||
| Llama-2 7B-chat | 0.797 | 0.586 | 0.661 | 0.595 | 0 | 0.703 | 0.533 | 0.648 | 0.499 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.788 | 0.555 | 0.644 | 0.558 | 0 | 0.707 | 0.537 | 0.596 | 0.502 | 0 | ||
| Flan-T5 XL | 0.757 | 0.517 | 0.515 | 0.511 | 0 | 0.678 | 0.521 | 0.587 | 0.489 | 0 | ||
| Phi-2 | 0.747 | 0.524 | 0.552 | 0.518 | 0 | 0.710 | 0.541 | 0.611 | 0.520 | 0 | ||
@@
In this table, ``IND" and ``OOD" indicates in-domain evaluation and out-of-domain evaluation, respectively; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric ``M-Rec", ``M-Pre", and ``M-F1" represents macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.

### v2

@@ C.2 Instructions on Product Relation Prediction ()
C.3 Instructions on Product Matching ()
@@
C.4 Instructions on Sentiment Analysis ()
Seed Instruction Given the userβs review, identify the userβs sentiment from the listed options. Answer using one of the options.
Generated Instruction 1 Assess the userβs sentiment in the provided review and select the appropriate sentiment option from the list as the answer.
@@ C.4 Instructions on Sentiment Analysis ()
Generated Instruction 3 Carefully assess the userβs review for any strong expressions of sentiment, either positive or negative. Based on your analysis, select the most fitting sentiment option from the provided list as output.
Generated Instruction 4 Analyze the userβs review text and determine the overall sentiment expressed, then choose the corresponding sentiment option from the provided list (A: very positive, B: positive, C: neutral, D: negative, E: very negative) based on the identified sentiment.
Unseen Instruction Analyze the userβs review and determine the sentiment based on the listed options.
C.5 Instructions on Sequential Recommendation ()
@@
Generated Instruction 2 Rank the items in options and predict the userβs next purchase from the listed options by analyzing her historical purchases.
Generated Instruction 3 The userβs purchase history implies her preferences. Rank the items in the options based on the userβs preferences. Output the item that the user is most likely to purchase next from the options.
@@
Unseen Instruction Estimate the userβs intent based on the userβs purchase history, and predict the next product that the user is most likely to purchase from the given options.
C.6 Instructions on Multi-class Product Classification ()
@@
C.7 Instructions on Product Substitute Identification ()
@@
Generated Instruction 3 Check if a product can function as a substitute for a given query, even if it doesnβt fully meet all requirements. Output yes if it can or no otherwise.
@@
C.8 Instructions on Query Product Ranking ()
@@
Unseen Instruction Evaluate the query against each productβs title, determine the relevance between the query and the product, and organize the products in descending order of relevance, ensuring that the product with the highest relevance is positioned at the top of the list.
C.9 Instructions on Answerability Prediction ()
@@
C.10 Instructions on Answer Generation ()
@@
Appendix D Training and Evaluation Details
The training details of general-purpose and e-commerce baseline LLMs are shown in TableΒ A5. The general-purpose LLMs undergo the 1-shot evaluation, which measures the sample with one in-context example and one test case. For each task, the 1-shot evaluation dataset is composed of all 1K test samples and the same amount of training samples (except for , which uses all 253 test samples and 253 training samples) from its own test and training sets. The e-commerce LLM undergoes both 1-shot and 0-shot evaluation.
| General-purpose LLMs | URL | Accessibility |
| GPT-4 Turbo | https://platform.openai.com/ | API |
| Gemini Pro | https://ai.google.dev/tutorials/python_quickstart | API |
| Claude 2.1 | https://docs.anthropic.com/claude/reference/getting-started-with-the-api | API |
| Llama-2 13B-chat | https://huggingface.co/meta-llama/Llama-2-13b-chat-hf | Checkpoint |
| Mistral-7B-Instruct-v0.2 | https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 | Checkpoint |
| EcomGPT | https://github.com/Alibaba-NLP/EcomGPT | Checkpoint |
| Tasks | Baseline Models | Model Source | Mode | Parameters |
| SUOpenTagΒ (Xu etΒ al., 2019) | https://github.com/hackerxiaobai/OpenTag_2019 | training | epoch: 100 | |
| AVEQAΒ (Wang etΒ al., 2020) | https://github.com/Zinc-30/aveqa | training | epoch: 89; batch size: 16 | |
| RGCNΒ (Schlichtkrull etΒ al., 2018) | https://github.com/JinheonBaek/RGCN | training | epochs: 500K | |
| DeBERTaV3Β (He etΒ al., 2022) | https://huggingface.co/microsoft/deberta-v3-base | fine-tuning | epoch: 3; Β Β Β Β batch size: 8 | |
| BERTweetΒ (Nguyen etΒ al., 2020) | https://huggingface.co/finiteautomata/bertweet-base-sentiment-analysis | training | epoch: 3; Β Β Β Β batch size: 8 | |
| DeBERTaV3Β (He etΒ al., 2022) | https://huggingface.co/microsoft/deberta-v3-base | fine-tuning | epoch: 3; Β Β Β Β batch size: 8 | |
| P5Β (Geng etΒ al., 2022) | https://github.com/jeykigung/P5 | evaluate on checkpoint | checkpoint: toy base | |
| gSASRecΒ (Petrov & Macdonald, 2023) | https://github.com/asash/gSASRec-pytorch | training | sequence length: 50 | |
| RecformerΒ (Li etΒ al., 2023) | https://github.com/AaronHeee/RecFormer | fine-tuning | epoch: 16; Β Β Β Β batch size 8 | |
| GPT-4 TurboΒ (OpenAI, 2023) | https://platform.openai.com/ | 1-shot Β Β Β learning | prompt template inΒ E.2 | |
| , , , , | BERTΒ (Devlin etΒ al., 2019) | https://huggingface.co/bert-base-multilingual-cased | fine-tuning | epoch: 3; Β Β Β Β batch size: 8 |
| DeBERTaV3Β (He etΒ al., 2022) | https://huggingface.co/microsoft/deberta-v3-base | fine-tuning | epoch: 3; Β Β Β Β batch size: 8 |
@@ Appendix D Training and Evaluation Details
In this table, βtrainingβ means training from scratch, βevaluate on checkpointβ means that we evaluate using the checkpoint provided by the link, βfine-tuningβ means that we fine-tune the checkpoint provided by the link on specific tasks, and β1-shot learningβ indicates that we directly use the model checkpoint from the link, and prompt the model with one in-context example.
The training details of SoTA task-specific models are presented in Table A6. For each task, we train and test on the SoTA task-specific models of each task using its own training, validation, and IND test sets (i.e., task-specific). For the tasks with the OOD test set, we save the trained model and test them on the OOD test set. For SUOpenTagΒ (Xu etΒ al., 2019) and AVEQAΒ (Wang etΒ al., 2020), we evaluate the task according to the equationΒ 1. Note that we do not consider MAVEQAΒ (Yang etΒ al., 2022) as the baseline of , as it has demonstrated similar performance to AVEQA. For RGCNΒ (Schlichtkrull etΒ al., 2018), we construct a product-product graph, in which the nodes are products, and the edges are specified by the relations between products (e.g., similar products). We initialize the node embeddings in the RGCN using the embeddings of product titles generated from BERTΒ (Devlin etΒ al., 2019). The BERT, DeBERTa, and BERTweet baselines are implemented through sentence-transformerΒ (Reimers & Gurevych, 2019). For P5Β (Geng etΒ al., 2022), we conduct zero-shot evaluations on beauty, sports, and toys base checkpoints, and report the best result (toy base). Similar to the RGCN, we use the BERT-generated embeddings of product titles to initialize item embeddings in the gSASRec. For both gSASRec and RecformerΒ (Li etΒ al., 2023), we evaluate the results of the next product of interest within the candidate list as detailed in AppendixΒ A.2.2
Appendix E Prompt Templates
The following prompts are used for evaluating general-purpose and e-commerce LLMs. For Claude 2.1, the prompt begins with {HUMAN PROMPT} and ends with {AI PROMPT}. For Mistral-7B Instruct-v0.2, the prompt is wrapped with β[INST]β.
E.1 Prompt Templates for Zero-shot Evaluation
@@
E.2 Prompt Templates for One-shot Evaluation
@@
Appendix F Analysis on Base Models
@@
In this table, βBase Modelβ presents the base models used in models. On each task, the best performance of , , and when using different base models is in bold.
@@
We compare the 6 base models considered for the series of models (i.e., , , and ) and show the performance comparison in TableΒ A7. From TableΒ A7, we observe that Llama-2 13B-chat is the best-performing base model for . The instruction-tuned Llama-2 13B-chat model demonstrates considerable improvement compared to instruction-tuned Flan-T5 XXL on 7 out of the 10 tasks. We also observe that Mistral-7B Instruct-v0.2 and Phi-2 are the best-performing base models for and , respectively. Particularly, with instruction tuning, Mistral-7B Instruct-v0.2 achieves a notable average improvement of 2.2% over Llama-2 7B-chat across the 10 tasks. Similarly, Instruction-tuned Phi-2 also outperforms Flan-T5 XL on 6 out of the 10 tasks and achieves similar performance with Flan-T5 XL on the rest 4 tasks. The variations of trainable size and focused aspect contribute to the distinct inherent capabilities of the base models, which play a crucial role in adapting LLMs to e-commerce scenarios.
Appendix G Complete Experimental Results
@@
The complete results for both IND and OOD evaluations for tasks (1) attribute value extraction (), (2) product relation prediction (), (3) product matching (), (4) sentiment analysis (), (5) sequential recommendation (), (6) multi-class product classification (), (7) product substitute identification (), (8) query product ranking (), (9) answerability prediction (), and (10) answer generation () are presented in TableΒ A8, Β A9, Β A10, Β A11, Β A16, Β A12, Β A13, Β A16, Β A14, andΒ A15, respectively. Overall, these tables show that models fine-tuned on outperform the general-purpose LLMs and SoTA task-specific models in the IND test. Meanwhile, the exhibits good generalizability to the OOD data. Because of the high-quality , achieves remarkable performance with different base models. Note that #failed in tables represents the number of failure cases for which we cannot extract meaningful results from the model output. These failure cases are counted as wrong predictions when calculating the metrics. NPR refers to the negative prediction rate.
Regarding the hallucination issue, we observe very limited hallucination from . As shown in the #failed column of TableΒ A8 and Β A16, in and tasks, could slightly suffer from hallucination and does not output results in the desired format for several testing samples. For the other tasks except for the task, as shown in TableΒ A9-A14, is robust to hallucination and could always output options in a desired format. For the task, we randomly sample 50 answers generated by and do not observe hallucination with manual checks.
| Model | IND | OOD | |||||||||
| Recall* | Precision* | F1* | #failed | Recall* | Precision* | F1* | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.422 | 0.598 | 0.495 | 6 | 0.317 | 0.529 | 0.397 | 1 | ||
| Gemini Pro | 0.318 | 0.523 | 0.396 | 4 | 0.203 | 0.426 | 0.275 | 6 | |||
| Claude 2.1 | 0.310 | 0.494 | 0.381 | 59 | 0.312 | 0.600 | 0.410 | 66 | |||
| Llama-2 13B-chat | 0.002 | 0.002 | 0.002 | 0 | 0.000 | 0.000 | 0.000 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.321 | 0.435 | 0.369 | 69 | 0.217 | 0.337 | 0.264 | 52 | |||
| E-commerce LLM | EcomGPT | 0.000 | 0.000 | 0.000 | 905 | 0.001 | 0.042 | 0.001 | 869 | ||
| SoTA task-specific model | SUOpenTag | 0.603 | 0.500 | 0.546 | 0 | 0.124 | 0.173 | 0.144 | 0 | ||
| AVEQA | 0.425 | 0.491 | 0.456 | 0 | 0.283 | 0.257 | 0.269 | 0 | |||
| Task-specific | Flan-T5 XXL | 0.298 | 0.519 | 0.378 | 7 | 0.362 | 0.701 | 0.477 | 0 | ||
| Llama-2 13B-chat | 0.544 | 0.666 | 0.599 | 3 | 0.448 | 0.613 | 0.518 | 2 | |||
| Llama-2 7B-chat | 0.531 | 0.660 | 0.588 | 1 | 0.323 | 0.499 | 0.392 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.720 | 0.799 | 0.757 | 5 | 0.374 | 0.544 | 0.443 | 0 | |||
| Flan-T5 XL | 0.258 | 0.449 | 0.328 | 7 | 0.276 | 0.538 | 0.365 | 3 | |||
| Phi-2 | 0.304 | 0.570 | 0.397 | 265 | 0.288 | 0.488 | 0.362 | 0 | |||
| Generalist | Flan-T5 XXL | 0.353 | 0.611 | 0.447 | 2 | 0.360 | 0.699 | 0.476 | 0 | ||
| Llama-2 13B-chat | 0.530 | 0.646 | 0.582 | 1 | 0.276 | 0.425 | 0.335 | 0 | |||
| Llama-2 7B-chat | 0.514 | 0.641 | 0.571 | 4 | 0.236 | 0.392 | 0.294 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.612 | 0.722 | 0.662 | 0 | 0.304 | 0.463 | 0.367 | 0 | |||
| Flan-T5 XL | 0.297 | 0.514 | 0.376 | 1 | 0.267 | 0.518 | 0.352 | 1 | |||
| Phi-2 | 0.455 | 0.578 | 0.509 | 0 | 0.237 | 0.417 | 0.302 | 0 | |||
@@
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together. Recall*, Precision* and F1* are defined as equationΒ 1 in AppendixΒ A.1.1, and #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. On each task, the best performance is in bold.
| Model | IND | OOD | |||||||||||
| Accuracy | M-Rec | M-Pre | M-F1 | #failed | Accuracy | M-Rec | M-Pre | M-F1 | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.384 | 0.487 | 0.381 | 0.326 | 0 | 0.488 | 0.496 | 0.392 | 0.392 | 0 | ||
| Gemini Pro | 0.128 | 0.385 | 0.352 | 0.136 | 1 | 0.147 | 0.359 | 0.390 | 0.123 | 0 | |||
| Claude 2.1 | 0.508 | 0.347 | 0.344 | 0.275 | 10 | 0.362 | 0.394 | 0.400 | 0.277 | 4 | |||
| Llama-2 13B-chat | 0.473 | 0.333 | 0.333 | 0.333 | 0 | 0.419 | 0.338 | 0.339 | 0.324 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.442 | 0.323 | 0.325 | 0.324 | 0 | 0.422 | 0.338 | 0.351 | 0.327 | 0 | |||
| E-commerce LLM | EcomGPT | 0.147 | 0.101 | 0.101 | 0.091 | 444 | 0.125 | 0.125 | 0.092 | 0.096 | 455 | ||
| SoTA task-specific model | DeBERTaV3 | 0.762 | 0.575 | 0.620 | 0.588 | 0 | 0.658 | 0.514 | 0.570 | 0.507 | 0 | ||
| RGCN | 0.615 | 0.665 | 0.637 | 0.506 | 0 | 0.576 | 0.373 | 0.372 | 0.356 | 0 | |||
| Task-specific | Flan-T5 XXL | 0.754 | 0.516 | 0.511 | 0.508 | 0 | 0.663 | 0.506 | 0.468 | 0.466 | 0 | ||
| Llama-2 13B-chat | 0.769 | 0.530 | 0.517 | 0.521 | 0 | 0.690 | 0.520 | 0.472 | 0.483 | 0 | |||
| Llama-2 7B-chat | 0.774 | 0.541 | 0.628 | 0.537 | 0 | 0.695 | 0.526 | 0.803 | 0.498 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.782 | 0.547 | 0.689 | 0.543 | 0 | 0.711 | 0.532 | 0.808 | 0.502 | 0 | |||
| Flan-T5 XL | 0.704 | 0.467 | 0.496 | 0.460 | 0 | 0.592 | 0.471 | 0.625 | 0.427 | 0 | |||
| Phi-2 | 0.584 | 0.372 | 0.379 | 0.348 | 0 | 0.406 | 0.349 | 0.334 | 0.251 | 0 | |||
| Generalist | Flan-T5 XXL | 0.769 | 0.531 | 0.517 | 0.522 | 0 | 0.703 | 0.533 | 0.648 | 0.499 | 0 | ||
| Llama-2 13B-chat | 0.775 | 0.599 | 0.635 | 0.611 | 0 | 0.726 | 0.564 | 0.611 | 0.558 | 0 | |||
| Llama-2 7B-chat | 0.797 | 0.586 | 0.661 | 0.595 | 0 | 0.703 | 0.533 | 0.648 | 0.499 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.788 | 0.555 | 0.644 | 0.558 | 0 | 0.707 | 0.537 | 0.596 | 0.502 | 0 | |||
| Flan-T5 XL | 0.757 | 0.517 | 0.515 | 0.511 | 0 | 0.678 | 0.521 | 0.587 | 0.489 | 0 | |||
| Phi-2 | 0.747 | 0.524 | 0.552 | 0.518 | 0 | 0.710 | 0.541 | 0.611 | 0.520 | 0 | |||
@@
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric βM-Recβ, βM-Preβ, and βM-F1β represents macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.